# Multi-Agent DRL for Multi-Objective Twin Migration Routing with Workload Prediction in 6G-enabled IoV

### v1

@@ I-A Background and Motivations
The Internet of Vehicles (IoV) enables network connections through vehicle-to-vehicle and vehicle-to-infrastructure communication, with mobile vehicles serving as information perception objects [1]. To enhance traffic operation efficiency, vehicles need to acquire real-time information from their surrounding environments and infrastructure through built-in wireless sensors. The Sixth-Generation (6G) wireless communication technology is anticipated to achieve ultra-reliable and low-latency communications with large-scale coverage [2], supporting intelligent applications while meeting the reliability and security demands of future communication networks. With the adoption of 6G technology, the coverage and real-time performance of IoV will be significantly improved, encompassing connections across space, air, and ground [3]. In 6G-enabled IoV networks, edge servers, such as RoadSide Units (RSUs), Unmanned Aerial Vehicles (UAVs), and satellites, can deploy computationally intensive vehicular applications, including 3D virtual human remote projection, augmented reality navigation, and autonomous driving [4], providing users with immersive and high-quality vehicular services [5].
@@
Vehicle Twins (VTs), functioning as an integrated Digital Twin (DT) technology, are high-precision virtual replicas of vehicles that combine physical vehicle models, sensing data, and historical data [6]. VTs play a crucial role in 6G-enabled IoV by collecting and analyzing traffic data to monitor vehicle states and predict traffic conditions [7]. However, constructing and updating VTs requires substantial computing and storage resources for real-time vehicle state monitoring and immersive service enhancement [8]. To address this, resource-constrained vehicles can offload VTs in the form of data packets to nearby edge servers (e.g., RSUs, UAVs, and satellites) after acquiring real-time data about themselves and their surroundings through multi-modal sensors, thus enabling more effective monitoring and simulation of vehicle states. Furthermore, with the support of 6G, vehicles can achieve reliable and ultra-low latency communications, ensuring real-time data updates and enhancing the responsiveness of vehicular applications.
@@
Due to the dynamic mobility of vehicles and the uneven distribution of edge servers, the service range of a single edge server cannot fully cover the trajectory of vehicles in practical application scenarios. Thus, it is challenging to ensure long-term stable communication interaction between a moving vehicle and its static VTs. To maintain the continuity of high-quality vehicular applications and services, VTs need to be dynamically migrated among edge servers based on the driving route of the corresponding vehicle. However, implementing effective VT migration presents several challenges:
@@
  • Challenges I. Nonuniform deployment of ground edge servers: The distribution of ground edge servers (i.e., RSUs) varies significantly across different regions. For example, in cities, RSUs are densely deployed and generally follow a more regular distribution pattern. In contrast, rural areas generally have a sparse and scattered distribution of edge servers, which may disrupt the continuity of VT migration.
@@ I-B Solutions and Contributions
To address these challenges, we propose a novel workload prediction-based VT migration framework in 6G-enabled IoV. Specifically, we introduce UAVs, leveraging their high mobility to complement RSU computing resources on busy roads during peak hours, and incorporate low-Earth orbit satellites to provide vehicular services in remote areas. To prevent VT migration interruptions during peak traffic hours, we propose a Long Short-Term Memory (LSTM)-based Transformer model to accurately predict the long-term workload of edge servers, thus avoiding VT migration to overloaded edge servers. Furthermore, considering the complex environment of 6G-enabled IoV involving simultaneous multiple VT migrations, it is difficult for traditional DRL algorithms to efficiently identify optimal VT migration routes. To overcome this limitation, we propose a Dynamic Mask Multi-Agent Proximal Policy Optimization (DM-MAPPO) algorithm. The DM-MAPPO algorithm dynamically determines whether all output actions meet the requirements and reduces the probability of unreachable actions to zero. Finally, the action sequence obtained by DM-MAPPO represents the most efficient VT migration route[10]. Our main contributions are summarized as follows:
@@
  • To efficiently obtain multiple optimal VT migration routes in the complex environment of 6G-enabled IoV networks, we propose a DM-MAPPO algorithm. Specifically, we integrate a dynamic mask module into the MAPPO algorithm, which sets the output probability of each invalid action to 00, thereby dynamically reducing the action space and enhancing the efficiency of identifying optimal VT migration routes (For Challenge III).
@@ II-B Workload Prediction of Edge Servers
To prevent VTs from being migrated to overloaded edge servers, it is essential to predict the long-term workload of edge servers in advance. In the field of time series prediction, the emergence of Recurrent Neural Networks (RNNs) has shown promise for long-term traffic flow forecasting [18]. Gated Recurrent Unit (GRU) and LSTM networks retain long-term memory through gating mechanisms and cell states, offering both effectiveness and lightweight properties [19, 20]. Based on the combination of graph neural networks and RNNs, the authors in [21] proposed a dynamic graph convolutional cycle network to improve traffic prediction accuracy by capturing dynamic correlation features between vehicle locations in the road network. In recent years, more prediction algorithms have been improved on the basis of Transformers and carried out long-term time series prediction[22]. For instance, the authors in [23] proposed the Pyraformer algorithm, which captures multi-resolution features through an inter-scale tree structure. In [24], the authors proposed the Informer algorithm, which utilizes a prob-sparse self-attention mechanism with a distillation technique to efficiently extract key information. Due to the strong periodicity of traffic flow data, these algorithms generally focus on capturing long-term periodicity information while overlooking some short-term time series variations. However, edge servers are highly sensitive to fluctuations in traffic flow data, making accurate workload prediction crucial for efficient VT migration decisions.
@@ III-A Network Model
In time slot $t$, the latitude and longitude of vehicle $v$ are denoted as $\phi_{v}^{t}$ and $\lambda_{v}^{t}$, respectively, while the latitude and longitude of edge server $m$ are represented as $\phi_{m}^{t}$ and $\lambda_{m}^{t}$, respectively. We denote the latitude difference between vehicle $v$ and edge server $m$ as $\Delta\phi_{v,m}^{t}=\frac{\pi}{180}(\phi_{v}^{t}-\phi_{m}^{t})$, and the longitude difference as $\Delta\lambda_{v,m}^{t}=\frac{\pi}{180}(\lambda_{v}^{t}-\lambda_{m}^{t})$. Based on the Haversine function, the central angular radian $\eta_{v,m}^{t}$ corresponding to the distance $D_{v,m}(t)$ between vehicle $v$ and edge server $m$ can be calculated as
@@
where the Haversine function is $\mathrm{hav}(\eta_{v,m}^{t})=\sin^{2}\Big{(}\frac{\eta_{v,m}^{t}}{2}\Big{)}=\sin^{2}(\Delta\lambda_{v,m}^{t})+\cos(\frac{\pi}{180}\phi_{v}^{t})\cos(\frac{\pi}{180}\phi_{m}^{t})\sin^{2}(\Delta\phi_{v,m}^{t})$.
@@
Satellites are generally considered to be in geosynchronous orbit at $35,786\>\rm{km}$ above the Earth’s surface, while low-Earth-orbit satellites typically operate at an altitude of $550\>\rm{km}$. In each time slot, the size of the data generated by vehicle $v$, including VT migration requests, is denoted as $S_{v}$. Based on the uplink transmission speed, the uplink latency $T_{v,m}^{up}(t)$ is given by
@@ III-B Computing Model
We define the size of pending data in the cache of the edge server $m$ in time slot $t$ as $S_{m}(t)$. Thus, the load of edge server $m$ in time slot $t$ can be calculated as
@@
$g_{v,k}$ =$=\begin{cases}0,&k\neq m^{t},\\ 1,&k=m^{t},\end{cases}$ (8b)
@@ III-C Migration Model
If vehicle $v$ is out of the service range of the connected edge server $m$, or the connected edge server $m$ is overloaded, the corresponding VT migration will fail. Thus, whether the VT migration of vehicle $v$ fails in time slot $t$ is given by
@@
$q_{v}=\frac{\sum_{t=0}^{t_{v}^{end}}{\xi_{v}^{t}}}{t_{v}^{end}}\times\text{100\%}.$% . (15)
@@ IV Problem Formulation
In this paper, we aim to optimize three objectives. The first goal is to minimize the total latency $T_{v,m}^{sum}(t)$ of all vehicles, thus ensuring a fully immersive experience. Secondly, the uneven distribution of RSUs in a region may cause cyclical and regional load peaks of edge servers. To balance the workload of all edge servers in the region, we take the variance of the workload of all edge servers $\mathrm{Var}(S_{m}^{\prime}(t))$ as one of the optimization objectives, and $\mathrm{Var}(S_{m}^{\prime}(t))$ can be calculated as $\mathrm{Var}(S_{m}^{\prime}(t))=\sum_{i=1}^{M}{(S_{i}^{\prime}(t)-\mu_{i}(t))^{2}}$, where $\mu_{i}(t)$ represents the average of $S_{m}^{\prime}(t)$. Finally, frequent VT migrations not only lead to additional unnecessary computing overhead but also degrade the quality of vehicular services. Thus, the goal is to minimize the total number of VT migrations after optimization, and the three optimization objectives along with their corresponding constraints are formulated as
@@ V LSTM-based Transformer Model for Workload Predictions of Edge Servers
The workload of edge servers often exhibits periodic patterns [32]. LSTM typically performs well in predicting periodic data [33], but it performs poorly for aperiodic data, especially in capturing fine-grained workload variations. In contrast, the Transformer model can effectively capture the relationships among vectors, making it a key technique for generating accurate contextual semantics in natural language processing [34]. In this section, we propose the LSTM-based Transformer model to predict the workload of edge servers. The structure of the proposed model is shown in Fig. 2.
@@ VI-A POMDP for VT Migration Routing Tasks
  • Dynamic Mask Module: When VTs of vehicle $v$ migrated to edge server $m$ based on action $a_{v,m}^{t}$ lead to $S_{m}^{\prime}(t)>L_{m}^{max}$ or $D_{v,m}(t)>D_{m}^{max}$ in time slot $t$, the action $a_{v,m}^{t}$ are infeasible. We define the infeasible action set as $\mathcal{M}_{v,t}^{infeasible}$. To prevent DRL agents from choosing an infeasible action, we utilize the dynamic mask module to set the probability of infeasible actions to 00 [26, 38], which accelerates agents to identify optimal migration routes. We define the mask list $\mathbf{K}_{v}^{t}=[K_{v,1}^{t},\ldots,K_{v,m}^{t},\ldots,K_{v,M}^{t}]$, where $K_{v,m}^{t}$ is given by
@@
[…]
    Figure 4 shows the complete calculation process of the dynamic mask module. After mapping the output of the LSTM-based Transformer prediction model to workload data, the dynamic mask module will add the task data stored in the cache to obtain the actual workload of edge server $m$ in time slot $t$. This is because the unfinished task in the last time slot $t-1$ will be stored in the cache. Finally, the dynamic mask module will determine if the edge server is about to be overloaded based on the size of the VT migration task. In addition, to ensure that VTs are not migrated to edge servers too far away, the dynamic mask module performs the same mask operation on edge servers beyond the maximum service range $D_{m}^{max}$.
1 Initialize actor network parameters $\theta_{v}$, $\theta_{v}^{old}$, and critic network parameters $w_{v}$;
2 Initialize environment $Env$ and replay buffer $Buf$;
@@
5       Store predicted workloads into $Buf$.
@@
8                   Get the observation $\mathbf{\widetilde{o}}_{v}^{t}$ from $Env$.
@@
12                   Get the reward $r_{v}^{t}$ and the next observation $\mathbf{\widetilde{o}}_{v}^{t+1}$ from $Env$.
13                   Store $\{\mathbf{o}_{v}^{t},\mathbf{o}_{v}^{t+1},a_{v}^{t},r_{v}^{t}\}$ into $Buf$.
@@
16            Get the reward list $\left\{r_{v}^{t}\right\}_{t=0}^{t_{v}^{end}}$ and the next observation list $\left\{\mathbf{o}_{v}^{t+1}\right\}_{t=0}^{t_{v}^{end}}$ from $Buf$.
17             Calculate the next value list $\left\{Q_{w_{v}}(\mathbf{o}_{v}^{t+1},a_{v}^{t+1})\right\}_{t=0}^{t_{v}^{end}}$ as ${target}$.
18             Calculate the value list $\left\{Q_{w_{v}}(\mathbf{o}_{v}^{t},a_{v}^{t})\right\}_{t=0}^{t_{v}^{end}}$ as ${value}$.
@@
26                   Calculate ${ratio}$ according to Eq. (28).

### v2

# arXiv:2505.07290v2[cs.NI] 01 Mar 2026

@@ I-A Background and Motivations
The Internet of Vehicles (IoV) enables network connections through vehicle-to-vehicle and vehicle-to-infrastructure communication, with mobile vehicles serving as information perception objects [52, 45]. To enhance traffic operation efficiency, vehicles need to acquire real-time information from their surrounding environments and infrastructure through built-in wireless sensors. The Sixth-Generation (6G) wireless communication technology is anticipated to achieve ultra-reliable and low-latency communications with large-scale coverage [28], supporting intelligent applications while meeting the reliability and security demands of future communication networks. With the adoption of 6G technology, the coverage and real-time performance of IoV will be significantly improved, encompassing connections across space, air, and ground [13]. In 6G-enabled IoV networks, edge servers, such as RoadSide Units (RSUs), Unmanned Aerial Vehicles (UAVs), and satellites, can deploy computationally intensive vehicular applications, including 3D virtual human remote projection, augmented reality navigation, and autonomous driving [36], providing users with immersive and high-quality vehicular services [40].
@@
Vehicle Twins (VTs), functioning as an integrated Digital Twin (DT) technology, are high-precision virtual replicas of vehicles that combine physical vehicle models, sensing data, and historical data [19]. VTs play a crucial role in 6G-enabled IoV by collecting and analyzing traffic data to monitor vehicle states and predict traffic conditions [9]. However, constructing and updating VTs requires substantial computing and storage resources for real-time vehicle state monitoring and immersive service enhancement [5], and uploading directly to cloud servers will introduce additional communication latency. To address this, edge computing technology is introduced. Vehicles can offload VTs in the form of data packets to nearby edge servers (e.g., RSUs, UAVs, and satellites) after acquiring real-time data about themselves and their surroundings through multi-modal sensors. Then, vehicles with limited computing resources and energy reserves can be more effectively monitored for their status. Furthermore, with the support of 6G, vehicles can achieve reliable and ultra-low latency communications, ensuring real-time data updates and enhancing the responsiveness of vehicular applications.
@@
Due to the dynamic mobility of vehicles and the uneven distribution of edge servers, the service range of a single edge server cannot fully cover the trajectory of vehicles in practical application scenarios, necessitating collaboration among multiple servers. Thus, it is challenging to ensure long-term stable communication interaction between a moving vehicle and its static VTs. To maintain the continuity of high-quality vehicular applications and services, VTs need to be dynamically migrated among edge servers based on the driving route of the corresponding vehicle. However, implementing effective VT migration presents several challenges:
@@
  • Challenge I. Nonuniform deployment of ground edge servers: The distribution of ground edge servers (i.e., RSUs) varies significantly across different regions. For example, in cities, RSUs are densely deployed and generally follow a more regular distribution pattern. In contrast, rural areas generally have a sparse and scattered distribution of edge servers, which may disrupt the continuity of VT migration.
@@ I-B Solutions and Contributions
To address these challenges, we propose a novel workload prediction-based VT migration framework in 6G-enabled IoV. Specifically, we introduce UAVs, leveraging their high mobility to complement RSU computing resources on busy roads during peak hours, and incorporate Low-Earth-Orbit (LEO) satellites to provide vehicular services in remote areas. When terrestrial networks are overloaded, compromised, or unavailable, satellite communication can serve as a critical emergency solution, ensuring the continuity of core VT services [4]. To prevent VT migration interruptions during peak traffic hours, we propose a Long Short-Term Memory (LSTM)-based Transformer model to accurately predict the long-term workload of edge servers, thus avoiding VT migration to overloaded edge servers. Furthermore, considering the complex environment of 6G-enabled IoV involving simultaneous multiple VT migrations, it is difficult for traditional DRL algorithms to efficiently identify optimal VT migration routes. To overcome this limitation, we propose a Dynamic Mask Multi-Agent Proximal Policy Optimization (DM-MAPPO) algorithm. The DM-MAPPO algorithm dynamically determines whether all output actions meet the requirements and reduces the probability of unreachable actions to zero. Finally, the action sequence obtained by DM-MAPPO represents the most efficient VT migration route[6]. The main contributions of this paper are summarized as follows:
@@
  • To efficiently obtain multiple optimal VT migration routes in the complex environment of 6G-enabled IoV networks, we propose a DM-MAPPO algorithm. Specifically, we integrate a dynamic mask module into the MAPPO algorithm, which sets the output probability of each invalid action to 0, thereby dynamically reducing the action space and enhancing the efficiency of identifying optimal VT migration routes (For Challenge III).
@@ II-B Workload Prediction of Edge Servers
To prevent VTs from being migrated to overloaded edge servers, it is essential to predict the long-term workload of edge servers in advance. In the field of time series prediction, particularly in traffic flow forecasting, traditional algorithms often struggle to deliver superior performance due to the complex temporal dependencies, high volatility, and nonlinear characteristics of traffic flow data. Fortunately, the emergence of Recurrent Neural Networks (RNNs) has made precise prediction of long-term time series data possible [7]. Gated Recurrent Unit (GRU) and LSTM networks retain long-term memory through gating mechanisms and cell states, offering both effectiveness and lightweight properties [2, 33]. Based on the combination of graph neural networks and RNNs, the authors in [21] proposed a dynamic graph convolutional cycle network to improve traffic prediction accuracy by capturing dynamic correlation features between vehicle locations in the road network. In recent years, more prediction algorithms have been improved on the basis of Transformers and carried out long-term time series prediction[50]. For instance, the authors in [22] proposed the Pyraformer algorithm, which captures multi-resolution features through an inter-scale tree structure. In [51], the authors proposed the Informer algorithm, which utilizes a prob-sparse self-attention mechanism with a distillation technique to efficiently extract key information. Since the strong periodicity of traffic flow data, these algorithms generally focus on capturing long-term periodicity information while neglecting some short-term time series variations. However, edge servers are highly sensitive to fluctuations in traffic flow data, making accurate workload prediction crucial for efficient VT migration decisions.
@@ III-A Network Model
In time slot $t$, the latitude and longitude of vehicle $v$ are denoted as $\phi_{v}^{t}$ and $\lambda_{v}^{t}$, respectively, while the latitude and longitude of edge server $m$ are represented as $\phi_{m}^{t}$ and $\lambda_{m}^{t}$, respectively. We denote the latitude difference between vehicle $v$ and edge server $m$ as $\Delta\phi_{v,m}^{t}=\frac{\pi}{180}(\phi_{v}^{t}-\phi_{m}^{t})$, and the longitude difference as $\Delta\lambda_{v,m}^{t}=\frac{\pi}{180}(\lambda_{v}^{t}-\lambda_{m}^{t})$. The conversion factor from degrees to radians $\vartheta$ is defined as $\vartheta=\frac{\pi}{180}$. Based on the Haversine function, the central angular radian $\eta_{v,m}^{t}$ corresponding to the distance $D_{v,m}(t)$ between vehicle $v$ and edge server $m$ can be calculated as
@@
where the Haversine function is $\mathrm{hav}(\eta_{v,m}^{t})=\sin^{2}\Big(\frac{\eta_{v,m}^{t}}{2}\Big)=\sin^{2}(\Delta\lambda_{v,m}^{t})+\cos(\vartheta\phi_{v}^{t})\cos(\vartheta\phi_{m}^{t})\sin^{2}(\Delta\phi_{v,m}^{t})$.
@@
Satellites are generally considered to be in geosynchronous orbit at $35,786\>\rm{km}$ above the Earth’s surface, while LEO satellites typically operate at an altitude of $550\>\rm{km}$. In each time slot, the size of the data generated by vehicle $v$, including VT migration requests, is denoted as $S_{v}$. Based on the uplink transmission speed, the uplink latency $T_{v,m}^{up}(t)$ is given by
@@ III-B Computing Model
Similar to RSUs, UAVs are considered to have certain computational capabilities [16], and we define the size of uncompleted tasks in the cache of the edge server $m$ in time slot $t$ as $S_{m}(t)$. Thus, the load of edge server $m$ in time slot $t$ can be calculated as
@@
$g_{v,k}$ $=\begin{cases}0,&k\neq m^{t},\\ 1,&k=m^{t},\end{cases}$ (8b)
@@ III-C Migration Model
If vehicle $v$ is out of the service range of the connected edge server $m$, or the connected edge server $m$ is overloaded, the VT service will be interrupted. Thus, whether the VT service of vehicle $v$ fails in time slot $t$ is given by
@@
$q_{v}=\frac{\sum_{t=0}^{t_{v}^{end}}{\xi_{v}^{t}}}{t_{v}^{end}}\times\text{100\%}.$ (15)
@@ IV Problem Formulation
In this paper, we aim to optimize three objectives. The first goal is to minimize the total latency $T_{v,m}^{sum}(t)$ of all vehicles, thus ensuring a fully immersive experience. Secondly, the uneven distribution of RSUs in a region may cause cyclical and regional load peaks of edge servers. To balance the workload of all edge servers in the region, we take the variance of the workload of all edge servers $\mathrm{Var}(S_{m}^{\prime}(t))$ as one of the optimization objectives, and $\mathrm{Var}(S_{m}^{\prime}(t))$ can be calculated as $\mathrm{Var}(S_{m}^{\prime}(t))=\frac{1}{M}\sum_{i=1}^{M}{(S_{i}^{\prime}(t)-\mu_{i}(t))^{2}}$, where $\mu_{i}(t)$ represents the average of $S_{m}^{\prime}(t)$. Finally, frequent VT migrations not only lead to additional unnecessary computing overhead but also degrade the quality of vehicular services. Thus, the goal is to minimize the total number of VT migrations after optimization, and the three optimization objectives along with their corresponding constraints are formulated as
@@ V LSTM-based Transformer Model for Workload Predictions of Edge Servers
The workload of edge servers, while exhibiting periodic patterns [43], presents significant prediction challenges due to its complex temporal dependencies, high volatility, and nonlinear characteristics. LSTM typically performs well in predicting periodic data [20], but it performs poorly for aperiodic data, especially in capturing fine-grained workload variations. In contrast, the Transformer model can effectively capture the relationships among vectors, making it a key technique for generating accurate contextual semantics in natural language processing [34]. In this section, we propose the LSTM-based Transformer model to predict the workload of edge servers. The structure of the proposed model is shown in Fig. 2.
@@ VI-A POMDP for VT Migration Routing Tasks
  • Dynamic Mask Module: When VTs of vehicle $v$ migrated to edge server $m$ based on action $a_{v,m}^{t}$ lead to $S_{m}^{\prime}(t)>L_{m}^{max}$ or $D_{v,m}(t)>D_{m}^{max}$ in time slot $t$, the action $a_{v,m}^{t}$ are infeasible. We define the infeasible action set as $\mathcal{M}_{v,t}^{infeasible}$. To prevent DRL agents from choosing an infeasible action, we utilize the dynamic mask module to set the probability of infeasible actions to 0 [17, 37], which accelerates agents to identify optimal migration routes. We define the mask list $\mathbf{K}_{v}^{t}=[K_{v,1}^{t},\ldots,K_{v,m}^{t},\ldots,K_{v,M}^{t}]$, where $K_{v,m}^{t}$ is given by
@@
    $a_{v}^{t}=\underset{m}{\mathrm{argmax}}(\mathbf{P}_{v}^{t}\mathbf{K}_{v}^{t}).$ (24)
    Figure 4 shows the complete calculation process of the dynamic mask module. After mapping the output of the LSTM-based Transformer prediction model to workload data, the dynamic mask module will add the task data stored in the cache to obtain the actual workload of edge server $m$ in time slot $t$. This is because the unfinished task in the last time slot t1t-1 will be stored in the cache. Finally, the dynamic mask module will determine if the edge server is about to be overloaded based on the size of the VT migration task. In addition, to ensure that VTs are not migrated to edge servers too far away, the dynamic mask module performs the same mask operation on edge servers beyond the maximum service range $D_{m}^{max}$.
1Initialize actor network parameters $\theta_{v}$, $\theta_{v}^{old}$, and critic network parameters $w_{v}$;
2 Initialize environment EnvEnv and replay buffer BufBuf;
@@
5   Store predicted workloads into BufBuf.
@@
8         Get the observation $\mathbf{\widetilde{o}}_{v}^{t}$ from EnvEnv.
@@
12         Get the reward $r_{v}^{t}$ and the next observation $\mathbf{\widetilde{o}}_{v}^{t+1}$ from EnvEnv.
13         Store $\{\mathbf{o}_{v}^{t},\mathbf{o}_{v}^{t+1},a_{v}^{t},r_{v}^{t}\}$ into BufBuf.
@@
16     Get the reward list $\left\{r_{v}^{t}\right\}_{t=0}^{t_{v}^{end}}$ and the next observation list $\left\{\mathbf{o}_{v}^{t+1}\right\}_{t=0}^{t_{v}^{end}}$ from BufBuf.
17      Calculate the next value list $\left\{Q_{w_{v}}(\mathbf{o}_{v}^{t+1},a_{v}^{t+1})\right\}_{t=0}^{t_{v}^{end}}$ as target{target}.
18      Calculate the value list $\left\{Q_{w_{v}}(\mathbf{o}_{v}^{t},a_{v}^{t})\right\}_{t=0}^{t_{v}^{end}}$ as value{value}.
@@
26         Calculate ratio{ratio} according to Eq. (28).